# Primordial Black Hole Formation in Rastall Gravity: Shifted Collapse Threshold and Exponential Abundance Sensitivity

### v1

@@ Abstract
Primordial black holes formed in the early universe are compelling candidates for dark matter. We investigate their production in Rastall gravity, a modification of general relativity that introduces a non-minimal coupling between matter and geometry through the non-conservation of the energy-momentum tensor. Analyzing cosmological perturbations during radiation domination, we demonstrate that the Rastall parameter fundamentally alters the collapse dynamics, modifying the growth of density fluctuations, the critical threshold for black hole formation, and the fluctuation amplitude at horizon crossing. Current cosmological constraints from Big Bang Nucleosynthesis, the cosmic microwave background, and large-scale structure restrict the Rastall parameter to small values, yet within this allowed range PBH production can be altered by orders of magnitude compared to general relativity. Our results establish primordial black holes as novel probes of modified gravity.
@@ 1 Introduction
Primordial black holes (PBHs) have emerged as compelling dark matter candidates and unique probes of the early universe [1, 2, 3, 4, 5, 6]. Formed from the gravitational collapse of large-amplitude density perturbations during radiation domination, PBHs bridge the gap between early universe physics and observational astronomy, particularly through gravitational wave astronomy [7, 8, 9, 10, 11, 12, 13]. The standard formation mechanism in general relativity (GR) requires specific enhancements of the primordial power spectrum, often necessitating fine-tuned inflationary dynamics [14, 15, 16, 17, 18, 19, 20, 21, 22, 23].
@@ 2 Rastall Gravity: Fundamentals and Background Evolution
While the background dynamics for pure radiation are identical to GR, differences emerge at the perturbation level. This makes PBH formation during radiation domination an ideal probe of Rastall gravity, as any deviations from GR must arise from perturbations rather than background expansion.
@@ 4.1.2 Horizon crossing ($x=1$)
$\delta_{H}(k)=\frac{4}{9}(1+\alpha_{H}\lambda)\mathcal{P}_{\mathcal{R}}^{1/2}(k),$ (39)
where $\alpha_{H}\approx 0.3$ is a numerical coefficient derived from matching, and $\mathcal{P}_{\mathcal{R}}(k)$ is the primordial curvature power spectrum.
@@ 5.1.2 Approach 2: Spherical collapse model
where $\gamma$ is a dimensionless coefficient encoding the response of the nonlinear collapse dynamics to the Rastall modification.

### v2

@@ Abstract
Primordial black holes formed in the early universe are compelling candidates for dark matter. We investigate their production in Rastall gravity, a modification of general relativity that introduces a non-minimal coupling between matter and geometry through the non-conservation of the energy-momentum tensor. Analyzing cosmological perturbations during radiation domination, we demonstrate that the Rastall parameter fundamentally alters the collapse dynamics, modifying the growth of density fluctuations, the critical threshold for black hole formation, and the fluctuation amplitude at horizon crossing if we consider the nearly conformal plasma. Current cosmological constraints from Big Bang Nucleosynthesis, the cosmic microwave background, and large-scale structure restrict the Rastall parameter to small values, yet within this allowed range PBH production can be altered by orders of magnitude compared to general relativity. Our results establish primordial black holes as novel probes of modified gravity.
@@ 1 Introduction
Primordial black holes (PBHs) have emerged as compelling dark matter candidates and unique probes of the early universe [1, 2, 3, 4, 5, 8, 6, 7, 8]. Formed from the gravitational collapse of large-amplitude density perturbations during radiation domination, PBHs bridge the gap between early universe physics and observational astronomy, particularly through gravitational wave astronomy [9, 10, 11, 12, 13, 14, 15]. The standard formation mechanism in general relativity (GR) requires specific enhancements of the primordial power spectrum, often necessitating fine-tuned inflationary dynamics [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29].
@@ 2 Rastall Gravity: Fundamentals and Background Evolution
While the background dynamics for pure radiation are identical to GR, differences emerge at the perturbation level if we consider the deviation from perfectly conformal plasma. The details are given in 2.1. This makes PBH formation during radiation domination an ideal probe of Rastall gravity, as any deviations from GR must arise from perturbations rather than background expansion.
2.1 Activation of Rastall Corrections in a Nearly Conformal Plasma
In the strict radiation-dominated limit with equation of state $p=\rho/3$, the energy–momentum tensor is traceless,
$T\equiv T^{\mu}_{\ \mu}=-\rho+3p=0.$ (18)
In this limit, the modified conservation law of Rastall gravity,
$\nabla_{\mu}T^{\mu\nu}=\gamma\nabla^{\nu}R,$ (19)
reduces identically to the standard conservation equation, and the theory becomes dynamically equivalent to General Relativity. Consequently, in an exactly conformal radiation fluid, Rastall corrections would be absent both at background and linear perturbation levels.
However, the primordial plasma is not exactly conformal. Around the QCD epoch and at temperatures relevant for horizon re-entry of small-scale perturbations, quantum effects generate a small but non-vanishing trace of the energy–momentum tensor due to the QCD trace anomaly and deviations from ideal radiation behaviour.
To capture this effect in a model-independent manner, we parametrize the deviation from conformality as
$p=\left(\frac{1}{3}-\delta\right)\rho,\qquad\delta\ll 1,$ (20)
which implies
$T=-\rho+3p=-3\delta\rho.$ (21)
Thus, while the background expansion remains effectively radiation-dominated, the trace is small but non-zero,
$T\neq 0,$ (22)
and the Rastall modification becomes active in a controlled manner.
At linear order,
$\delta T=-\delta\rho+3\delta p=-3\delta\,\delta\rho,$ (23)
so that perturbatively the trace does not vanish. The modified conservation law therefore contributes at first order in $\delta$, leading to corrections proportional to the product $\gamma\delta$ in the scalar perturbation equation.
@@ 4.1.2 Horizon crossing ($x=1$)
$\delta_{H}(k)=\frac{4}{9}(1+\gamma\lambda)\mathcal{P}_{\mathcal{R}}^{1/2}(k),$ (45)
where $\gamma$ is a numerical coefficient derived from matching, and $\mathcal{P}_{\mathcal{R}}(k)$ is the primordial curvature power spectrum.
@@ 5.1.2 Approach 2: Spherical collapse model
where $\gamma$ is a dimensionless coefficient encoding the response of the nonlinear collapse dynamics to the Rastall modification. Equation 51 should therefore be interpreted as the leading-order perturbative evolution equation in a nearly conformal plasma, where the effective deviation from GR is controlled by the small parameter
$\gamma_{\rm eff}\equiv\gamma\delta.$ (52)